### v3

@@ 4 Experimental Results and Analysis
Due to the lack of unified evaluation and comprehensive analysis, we run existing open-source codes in image-based virtual try-on area and evaluate them with the same referees. In this section, experimental results are given as quantitative results, qualitative results and user study. Besides the observations towards these results, suggestions in terms of method design are also analyzed. Previous methods adopt VITON while recent methods use VITON-HD for validation, so we keep the original dataset adopted by representative methods for fair comparison and use two methods (i.e., HR-VITON [61] and VITON-HD [18]) as a bridge evaluated by both datasets.
Figure 8: Quantitative results on VITON-HD.
@@ 4.1 Quantitative Results
We consider two testing conditions to quantitatively evaluate current representative approaches, i.e., paired condition and unpaired condition. Ground truth exists for paired condition while for unpaired condition the clothes in clothing image and person image are different. Unpaired condition shows a practical scenario but this situation lacks ground truth. With the human parsing method SCHP [67], partial comparisons are proposed for the first time, which separately evaluate the performance of clothing warping (denoted as “clothing”) and generation of non-try-on area (denoted as “person”). The quantitative results of representative methods in the order of publication year are shown in Fig. 7 and 8. As unpaired condition lacks ground truth, it is hard to evaluate the generation quality of full body for SSIM, LPIPS and Semantic Score. From the results, we have the following observations:
Figure 9: Qualitative results on VITON. We show the visual results with variations of person poses, shapes and clothing patterns.
@@
  • For high-resolution methods, HR-VTON [61] (grey circle) performs better, which attributes to its dual-path pipeline that simultaneously predict human body and clothing and the deformed clothing is further aligned with the predicted semantic segmentation.
  • With VITON-HD and HR-VITON as a bridge, by comparing Fig. 7 and 8 we can find that the overall generation quality of Diffusion-based methods outperforms GAN-based methods.
@@ 4.2 Qualitative Results
The qualitative results are displayed in Fig. 9 and 10 by showing several cases of image-based virtual try-on. The $1^{st}$ row in both figures shows an easy task where these representative methods are good at, and we can find that recent methods (i.e., in Fig. 10) generate results with high-fidelity and fine-grained details. The preservation of clothing texture such as special patterns and stripes (e.g., the $2^{nd}$, $3^{rd}$ and $4^{th}$ rows) could be further improved. GAN-based methods usually generate blurry results while Diffusion-based methods face challenges in controllability. Person shape (e.g., $4^{th}$ row) has little effects while challenging pose (e.g., the last row) will increase the difficulty of try-on. Challenging cases will be further discussed in Sec. 5.
Figure 10: Qualitative results on VITON-HD. We show the visual results with variations of person poses, shapes and clothing patterns.
From qualitative results, Style-Flow-VTON [40] performs best under low resolution in Fig. 9. For earlier works such as VITON [37], SP-VTON [111] and CP-VTON [119] models, the focus is on the processing of the try-on part. When trying on simple clothes, the center area of the clothes has a better generation effect, but the clothes and person parts in the non-try-on area are blurred. LM-VTON [75] and CP-VTON+ [82] have improved the above problems to some extent by preserving the image contents in the non-try-on area, but also contain artifacts. Although DCTON [30] adopts a different model framework from the VITON series, its performance is close to that of LM-VTON and CP-VTON+. The quality of the results generated by PFAFN [31], Style-Flow-VTON [40] and SDAFN [5] gets better, where the boundary between the body area and the clothing area is more natural and clear. It mainly attributes to the warping ability of flow. VITON-HD [18] and HR-VITON [61] are two methods designed for high-resolution virtual try-on, and the results of HR-VITON contain more details.
Overall, the results of Diffusion-based methods (in Fig 10) show clear appearance. GP-VTON [126] is a GAN-based method, but shows competitive performance, which attributes to its clothing warping strategy of separately warping partitions. DCI-VTON [34] first warps clothing and then use diffusion model to refine the try-on results, which both controls the fidelity of original clothes and shows clear appearance. Although Diffusion-based methods (e.g., Ladi-VTON [84], CAT-DM [136] and Stable VITON [56]) show clear appearance, the controllability in clothing color and texture can still be improved.
@@ 4.3 User Study
In order to assess the visual quality of the virtual try-on results in human perspective, we conducted a survey in the form of a questionnaire to gather feedback from people. The questionnaire consisted of $22$ evaluations, where for each algorithm two random samples were selected. Participants were asked to choose one of the five options to rate the quality of the try-on results for each sample: A. 5 (Very realistic), B. 4 (With minor flaws), C. 3 (Lack of details), D. 2 (Contain obvious artifacts), and E. 1 (Messy). We conducted the testing publicly on the internet and received a total of $333$ valid responses from participants. As shown in Fig. 11, the length of the colored areas in the figure represents the proportion of the corresponding evaluation score. The results show that the try-on performance of existing image generation methods is far from satisfactory, which could also be observed in Fig. 9 that humans have high demands from the try-on procedure. It remains a challenging task to generate results with natural warping and clear patterns.
Figure 11: The user study results on the qualitative results. Participants evaluate the randomly selected try-on images. The length of the colored areas represents the proportion of the corresponding evaluation level where the satisfactory decreases from A to E.
@@ 4.4 Discussion on Method Design
  • Pipeline perspective. The structures of existing pipelines are classified as shown in Fig. 2, and the specific pipeline adopted by each approach is illustrated in Table 1. It could be found that there is no obvious pipeline preference in the development of image-based virtual try-on methods, and recent methods tend to design one-stage pipeline (i.e., type \@slowromancapi@) with large-scale models. Except recent diffusion-based methods, the dual-path pipeline (i.e., type \@slowromancapvii@) achieves superior performance where the dual paths facilitate each other to optimize the generation performance.
@@
Figure 13: Performance under challenging poses on VITON. $1^{st}$ row: the results in a classic arms-crossed posture. $2^{nd}$ row: a more challenging case in arms occlusion where the color of sleeves is the same as torso. $3^{rd}$ row: the fine-grained details of bent wrist cannot be well recovered. $4^{th}$ row: putting on a piece of long-sleeved clothes under the arms-crossed posture.
Figure 14: Performance under challenging poses on VITON-HD.
5 Unresolved Issues
As illustrated in Table 1, i.e., person representation, most methods rely on human parsing. Human parsing plays a vital role in image-based virtual try-on, but also causes the main issues due to imperfect parsing ability. In this section, we further show examples for issues caused by human parsing.
5.1 Challenging Poses
Challenging pose usually attributes to the occlusion caused by arms. Take the arm-crossed posture ($1^{st}$ row of Fig. 13) for example, three problems should be overcome to perform well: a. correctly judge the position of the arm in front or back of the body and generate the right mask to put on clothes; b. correctly generate the arms and hands in a crossed state; and c. if the try-on is a long-sleeved clothing, correctly handle the relationship between the clothing on the arm and the exposed skin.
For problem a, from Fig. 13 we can find that most methods can tell the front or back position correctly (e.g., the results in $1^{st}$ row) except the ambiguous clothing with plain color (e.g., the results of PFAFN [31] and SDAFN [5] in the $2^{nd}$ row). For problem b, the crossed arms are not well solved. HR-VTON [61] and VITON-HD [18] mitigate this problem by optimizing the misaligned parts of the human body and clothing within the generator.
Figure 15: Changed clothing styles (VITON). The styles of target clothes cannot be preserved due to introducing original clothing information in person representation or clothing bias in the training dataset.
The results in the $4^{th}$ row reflect the model’s ability of handling problem c, where the character aims to wear a piece of long-sleeved clothes. Most methods fail to correctly put on sleeves due to the limitation of clothing warping and try-on module. In addition, the arms-crossed pose contains the crossing of the fingers and the arm, which requires high granularity generation ability of the model. Take the $3^{rd}$ row for example, current representation such as Openpose [13], Densepose [35] and semantic segmentation map [33, 67] cannot label the fine-grained pose at the wrist (Fig. 12), which fundamentally leads to the poor generation quality.
With the advent of large-scale image generation models [129, 142, 16], the issue of blurred body has been significantly improved. As shown by the results of Diffusion-based methods in Fig. 14, the position of arms and details of figures are well perceived. However, it should be pointed out that current large-scale models are not fully controllable in terms of the generated contents.
Figure 16: Changed clothing styles (VITON-HD).
@@ 5.2 Limited Human Parsing
Besides the pose estimation, image-based virtual try-on methods also rely on some other pre-processing results (as aforementioned in Sec. 2.3). Fig. 17 shows some results generated by HR-VITON and VITON-HD using noisy semantic information, where the wrong parser area has artifacts. It could be found in Fig. 4 that $\mathcal{P}_{\begin{subarray}{c}6,7,10,11\end{subarray}}$ involve information of original clothes expected to take off. For methods which adopt such representation (e.g., VITON [37], CP-VTON [119] and LM-VTON [75]), the try-on results are affected by the style of original clothes. Take an example for these methods, the collar style of target clothes changes to that of the original clothes as shown in Fig. 15. Unlike these methods, approaches like Style-Flow-VTON [40], VITON-HD [18], and HR-VTON [61], which utilize $\mathcal{P}_{\begin{subarray}{c}3,12\end{subarray}}$ to address the issue of neckline boundary residuals by redrawing the neckline area, particularly shown in the $3^{rd}$ and $4^{th}$ rows of Fig. 15.
Clothing style is an important aspect to show the performance of virtual try-on, but the current models has poor performance on unconventional styles. From top to bottom in Fig. 15 are short T-shirts, loose-try-on tops, long T-shirts and light tops. Fig. 16 also shows some examples of unconventional styles in VITON-HD dataset. Compared with the correct wearing state of the corresponding clothing in the rightmost column of each row, the try-on results generated by the model have obvious biases, and the model tends to make the clothing fit perfectly on the upper body of the person body, presenting a slimming effect. This is unfair for clothing with loose styles or special designs in length, such as the results in the first and third rows of Fig. 15, where the short T-shirts are forcibly stretched to the normal length of T-shirts, and the long T-shirts are forcibly reduced/truncated to the normal length of T-shirts. Essentially, the model does not really understand the clothing to be tried on, and it only deforms the clothing image and fills it in the original clothing area for replacement. The main reason attributes to that the model is trained with paired data where target clothes and the clothes on person are actually the same one, and in reality it is difficult to collect unpaired data. Clothing-agnostic person representation still deserve exploration.
Figure 17: Artifacts caused by noisy parsing.
@@ 5.3 Limited Clothing Parsing
It is worth mention that clothing parsing is not necessary, e.g., for some large-scale image generation models, where only the character image and clothing image are needed as conditions, and a mask is used to guide the area to be tried on. Due to the strong feature parsing ability and scalability of large image generation models, the dependency on parsing will be also alleviated.
@@ 6 Future Work
Current visual results show that there is still significant room for improvement in image-based virtual try-on methods. On one hand, person representation does not totally get rid of the original clothes, which will affect the clothing style of target clothes. On the other hand, clothing warping approaches are not flexible enough, especially for challenging poses. The emergence of large-scale models alleviates this issue to some extent, but controllable generation becomes a new challenge. In this section, we first continue to show some potential improvements along current research directions in aspects of person representation, clothing deformation, controllable generation, datasets, loss functions and evaluation metrics. Once the quality of generation becomes satisfactory, additional ideas such as size-aware try-on, fashion customization, and scenario-specific generation and so on can also be taken into account.
@@
Clothing-agnostic person representation. As the training data is paired where the person already wears target clothes, it will harm the generalization ability if person representation contains clothes information. Current representations $\mathcal{P}_{\begin{subarray}{c}3,6,7,10,11\end{subarray}}$ in Fig. 4 contains clues from the original clothes such as clothing contour, which will affect the clothing style of target clothes. Totally deleting the body area (i.e., $\mathcal{P}_{\begin{subarray}{c}4,12\end{subarray}}$ will get rid of original clothes but lose person characteristics such as skin color. Constructing triplets where a person wears different clothes could be a solution, but current methods [47, 31, 40] still rely on parser-based methods to construct the triplet, which inherits the limitation of human parser. Future direction could try to construct training triplets with parser-free approaches such as recent large-scale models, or totally delete original clothes but keep a sampler at exposed skin. For diffusion models based on mask inpainting, how to decouple the mask with original clothes also deserves exploring.
@@
Natural clothing deformation. Transformation under intrinsic rules such as STN and TPS is limited for flexible deformation. For example, they cannot put sleeves before the torso. Flow estimation learns pixel offset which improves flexibility but cannot generates new pixels that are not displayed in target clothing image but are shown under current posture. Implicit transformation shows potentials in clothing deformation and deserves further exploration, especially driven by a 3D deformation model. Additionally, existing methods distort target clothing image into the try-on area and ignore the original style. For example, a cropped top is stretched or a long coat is shrunken to normal length. Fine-grained clothing landmarks could be helpful for this situation.
@@
Controllable generation with large-scale model. Large-scale image generation models show powerful generalization and high quality, which is appealing for virtual try-on tasks. However, current models only meet semantic-level control such as satisfying the description from text. More precise and strict control has not been well solved such as preserving the original styles (e.g., Fig. 16). Effectively encoding clothing image and injecting to the generation model, or just using large-scale models to refine the warped clothes deserve a try.
@@
Flexible loss functions. Current training datasets only contain paired samples where the person in image already wears target clothes, so existing loss functions are designed to constrain the full image. However, existing methods cannot generalize well under unpaired evaluation. It is expected to explore flexible loss functions to support unpaired training, which respectively constrains try-on area with target clothing image and non-try-on area from person image.
@@
Size-aware try-on. Existing methods distort target clothing image into the inpainting area with bias towards original clothes. On one hand, the size design of target clothes, e.g., close-fitting or oversized, is ignored. On the other hand, the size of body shapes has not been taken into account. In the future, relative relationship between target clothes and person could be given with several keypoints. How to utilize such information to facilitate the generation process deserves researching.
Tailorable try-on. If size-aware function has been realized, we can further add tailor option by controlling clothing landmarks. For example, we can shorten the length of clothes by moving related landmarks. Further mapping language to landmark movements will make this application more practical.
Fashion customization. With the development of large generation models, future try-on tasks can be extended to fashion customization such as changing clothing color, pattern and style. Adding extra patterns can also be easily implemented by putting on specific images.

### v4

@@ 4 Experimental Results and Analysis
Due to the lack of unified evaluation and comprehensive analysis, we run existing open-source codes in image-based virtual try-on area and evaluate them with the same referees. In this section, experimental results are given as quantitative results, qualitative results and user study. Besides the observations towards these results, suggestions in terms of method design are also analyzed.
@@ 4.1 Quantitative Results
Previous methods are quantitatively evaluated in terms of testing conditions and areas. Fig. 8 shows two testing conditions, i.e., paired and unpaired. Paired situation refers to that we have a pair of clothing image and image of person wearing the same clothes. Such data are abundant on the Internet and usually used for training by masking the person image to obtain clothing agnostic person representation. Unpaired situation is close to practical scenario where the clothing in the person image is different from target clothing, where we do not have the ground truth for try-on image.
Figure 9: Three testing conditions (“All”, “Person” and “Clothing”). With the help of a parser [70], we separate the generated image to non-try-on area and clothing area. The deleted areas are replaced with black pixel.
We for the first time separately evaluate the generation quality of different areas for virtual try-on, i.e., the warped clothing area, the non-try-on area, and the whole area, which are respectively denoted as “Clothing”, “Person” and “All”. Fig. 9 shows several examples for different evaluation areas. The parsing labels of ground truth via a recent parser [70] are used to extract specific area for comparison under paired situation. For unpaired case, we should separately parse the ground truth and generated results for comparison. It should be noted that the parsers in training and testing for the comparison methods keep the same as the official implementation, which might be different from the parser used in the separated evaluation. And the parser for evaluation could be changed with the advancement of human parsing.
Figure 10: Cross-dataset quantitative results on VITON-HD. The vertical axis indicates the testing condition is paired or unpaired and the evaluation target is clothing, non-try-on area (denoted as “person”) or all together. For method symbol, the cross symbol indicates the clothing warping strategy is flow while the triangle symbol represents TPS-based method. VITON-HD* and HR-VITON* denote that we resize the image resolution from $1024\times 768$ to $256\times 192$ via the default resize function in python imaging library.
Figure 11: Quantitative results on VITON-HD. VITON-HD* and HR-VITON* denote that we resize the image resolution from $1024\times 768$ to $512\times 384$ via the default resize function in python imaging library.
The quantitative results of representative methods in the order of publication year are shown in Fig. 10 and 11. As unpaired condition lacks ground truth, it is hard to evaluate the generation quality of full body for SSIM, LPIPS and Semantic Score. FID could be computed because it measures the similarity of two image sets. It should be noted that the evaluation criteria are sensitive to image resolution, so the quantitative results are compared under the same resolution. The image resolution of the comparison methods has been illustrated in Sec. 3.3. To save space, we did not draw separate figure for VITON-HD [17] and HR-VITON [62], which are under the resolution of $1024\times 768$. We put their results to both Fig. 10 and 11.
From the quantitative results, we have the following observations:
Figure 12: Qualitative results on VITON-HD, with variations of person poses, shapes and clothing patterns. The methods marked with * use the officially pretrained weights trained on VITON for inference.
@@
  • For high-resolution methods, HR-VTON [62] performs better in Fig 10, which attributes to its dual-path pipeline that simultaneously predict human body and clothing and the deformed clothing is further aligned with the predicted semantic segmentation.
  • With VITON-HD and HR-VITON as a bridge, by comparing Fig. 10 and 11 we can find that the overall generation quality of diffusion-based methods outperforms GAN-based methods.
@@ 4.2 Qualitative Results
The qualitative results are displayed in Fig. 12 by showing several cases of image-based virtual try-on. The $1^{st}$ row the figure shows an easy task where these representative methods are good at, and we can find that recent methods generate results with high-fidelity and fine-grained details. The preservation of clothing texture such as special patterns and stripes (e.g., the $2^{nd}$, $3^{rd}$ and $4^{th}$ rows) could be further improved. GAN-based methods usually generate blurry results while diffusion-based methods face challenges in controllability. Person shape (e.g., $4^{th}$ row) has less effects while challenging pose (e.g., the last row) will increase the difficulty of try-on. Challenging cases will be further discussed in Sec. 5.
For earlier works such as VITON [38], SP-VTON [115] and CP-VTON [123] models, the focus is on the processing of the try-on part, which roughly align the clothes to body. The quality of the results generated by PFAFN [32], Style-Flow-VTON [41] and SDAFN [5] gets better, where the boundary between the body area and the clothing area is more natural and clear. It mainly attributes to the warping ability of flow. VITON-HD [17] and HR-VITON [62] are two methods designed for high-resolution virtual try-on, and the results of HR-VITON contain more details. Overall, the results of diffusion-based methods (in Fig 12) show clear appearance. GP-VTON [130] is a GAN-based method, but shows competitive performance, which attributes to its clothing warping strategy of separately warping partitions. DCI-VTON [35] first warps clothing and then use diffusion model to refine the try-on results, which both controls the fidelity of original clothes and shows clear appearance. Although diffusion-based methods (e.g., Ladi-VTON [87], CAT-DM [139], Stable VITON [57] and TPD [136]) show clear appearance, the controllability in clothing color and texture can still be improved.
Figure 13: The interface of user study.
(a) Gender
(b) Age
(c) Frequency
Figure 14: Demography of participants: (a) gender, (b) age and (c) frequency of shopping online.
@@ 4.3 User Study
In order to assess the visual quality of virtual try-on results in human perspective, we collect public questionnaires online. The questionnaire is designed to evaluate user preference towards 16 comparison methods in aspects of clothing accuracy and fitting realism. Clothing accuracy aims to reflect the quality of clothing reconstruction and fitting realism depends on the alignment bewteen garment and human body.
Different from straightforwardly rating 16 results, we choose MaxDiff (Maximum Difference Scaling, also known as Best-Worst Scaling) for our evaluation. We choose MaxDiff for the following main reasons: Firstly, humans are very picky about virtual try-on results, which may give similar low ratings to existing methods (leading to bias or skewed results). MaxDiff forces participants to make distinct choices, which reduces the likelihood of neutral or invalid responses and provides clearer insights into preferences. Secondly, choosing the most and least preferred items is more intuitive and straightforward than rating each method on a scale, making the process less cognitively demanding and the result objective.
Figure 15: User preference towards existing methods.
As shown in the interface figure (Fig. 13), to ensure that participants have enough patience to answer, each questionnaire contains 4 samples and for each sample we set up 6 best-worst questions with 3 options. The 4 samples (2 for clothing accuracy and 2 for fitting realism) in each questionnaire are randomly selected, and different questionnaires are probably different due to the large pool and random setting.
Figure 16: Performance under challenging poses on VITON-HD. The methods marked with * use the officially pretrained weights trained on VITON for inference.
We received a total of $139$ valid questionnaires from anonymous participants, based on which we compute the demography of participants (Fig. 14) and user preference (Fig. 15). The score for user preference is calculated via MaxDiff based on multinomial logit model. From Fig. 15 we have the following two observations: 1) Current methods show similar performance in clothing accuracy and fitting realism; and 2) Earlier GAN-based methods could hardly satisfy users for virtual try-on and recent methods improve a lot, which can also be validated in the visual results.
@@ 4.4 Discussion on Method Design
  • Pipeline perspective. The structures of existing pipelines are classified as shown in Fig. 2, and the specific pipeline adopted by each approach is illustrated in Table 1. It could be found that there is no obvious pipeline preference in the development of image-based virtual try-on methods, and recent methods tend to design one-stage pipeline (i.e., type \@slowromancapi@) with diffusion models. Except recent diffusion-based methods, the dual-path pipeline (i.e., type \@slowromancapvii@) achieves superior performance where the dual paths facilitate each other to optimize the generation performance.
@@
5 Unresolved Issues
As illustrated in Table 1, most methods rely on human parsing. Human parsing plays a vital role in image-based virtual try-on, but also causes the main issues due to imperfect parsing ability. In this section, we further show examples for issues caused by human parsing.
5.1 Challenging Poses
Challenging pose usually attributes to the occlusion caused by arms. Take the arm-crossed posture (Fig. 16) for example, three problems should be overcome to perform well: a. correctly judge the position of the arm in front or back of the body and generate the right mask to put on clothes; b. correctly generate the arms and hands in a crossed state; and c. correctly handle the relationship between the clothing on the arm and the exposed skin.
@@ 5.1 Challenging Poses
For problem a, from Fig. 16 we can find that most methods can tell the front or back position correctly except the ambiguous clothing with plain color. For problem b, the crossed arms are not well solved. HR-VTON [62] and VITON-HD [17] mitigate this problem by optimizing the misaligned parts of the human body and clothing within the generator. The arms-crossed pose contains the crossing of the fingers and the arm, which requires high granularity generation ability of the model. Take Fig. 17 for example, current representation such as Openpose [13], Densepose [36] and semantic segmentation map [34, 70] cannot label the fine-grained pose at the wrist, which fundamentally leads to the poor generation quality. Recent diffusion-based methods show high realism at fine-grained appearance. The relationship between clothing and body is also well addressed with the development of methods. However, it should be pointed out that current diffusion models are not fully controllable in terms of the generated contents.
Figure 18: Artifacts caused by noisy parsing.
@@ 5.2 Limited Human Parsing
Figure 19: Changed clothing styles on VITON-HD. The right-most figures show the correct state of target clothing. The methods marked with * use the officially pretrained weights trained on VITON for inference.
Besides the pose estimation, image-based virtual try-on methods also rely on some other pre-processing results (as aforementioned in Sec. 2.3). Fig. 18 shows some results generated by HR-VITON and VITON-HD using noisy semantic information, where the wrong parser area has artifacts. It could be found in Fig. 4 that $\mathcal{P}_{\begin{subarray}{c}6,7,10,11\end{subarray}}$ involve information of original clothes expected to take off. For methods which adopt such representation (e.g., VITON [38], CP-VTON [123] and LM-VTON [78]), the try-on results are affected by the style of original clothes. Take an example for these methods, the collar style of target clothes changes to that of the original clothes as shown in Fig. 19.
Clothing style is an important aspect to show the performance of virtual try-on. Previous efforts have been made to adjust the style or size of garments via association with landmarks or skeleton [15, 132, 68]. However, as shown in Fig. 19, it is difficult to keep the original style when the target clothing has unconventional styles. Compared with the correct wearing state of the corresponding clothing in the rightmost column of each row, the try-on results generated by the model have obvious biases, and the model tends to make the clothing fit perfectly on the upper body of the person body, presenting a slimming effect. This is unfair for clothing with loose styles or special designs in length, such as the results in the second and third rows of Fig. 19, where the short T-shirts are forcibly stretched to the normal length of T-shirts, and the long T-shirts are forcibly reduced/truncated to the normal length of T-shirts.
Essentially, the model does not really understand the clothing to be tried on, and it only deforms the clothing image and fills it in the original clothing area for replacement. The reliance on the parsing of original clothes also introduces bias to inpainting area.
5.3 Limited Clothing Parsing
@@
It is worth mention that clothing parsing is not necessary, e.g., for recenet diffusion-based models, where only the character image and clothing image are needed as conditions, and a mask is used to guide the area to be tried on. Due to the strong feature parsing ability and scalability of large image generation models, the dependency on parsing will be also alleviated.
@@ 6 Future Work
Current visual results show that there is still significant room for improvement in image-based virtual try-on methods. On one hand, person representation does not totally get rid of the original clothes, which will affect the clothing style of target clothes. On the other hand, clothing warping approaches are not flexible enough, especially for challenging poses. The emergence of diffusion models alleviates this issue to some extent, but controllable generation becomes a new challenge. In this section, we first continue to show some potential improvements along current research directions in aspects of person representation, clothing deformation, controllable generation, acceleration, datasets and evaluation metrics. Once the quality of generation becomes satisfactory, additional ideas such as mask-free in-the-wild try-on and scenario-specific generation can also be taken into account.
@@
Clothing-agnostic person representation. As the training data is paired where the person already wears target clothes, it will harm the generalization ability if person representation contains clothes information. Current representations $\mathcal{P}_{\begin{subarray}{c}3,6,7,10,11\end{subarray}}$ in Fig. 4 contains clues from the original clothes such as clothing contour, which will affect the clothing style of target clothes. Totally deleting the body area (i.e., $\mathcal{P}_{\begin{subarray}{c}4,12\end{subarray}}$ will get rid of original clothes but lose person characteristics such as skin color. Constructing triplets where a person wears different clothes could be a solution, but current methods [48, 32, 41] still rely on parser-based methods to construct the triplet, which inherits the limitation of human parser. Unpaired try-on has been explored, but is still challenging facing the lack of supervised training. Future direction could try to construct training triplets with parser-free approaches such as recent diffusion models, or totally delete original clothes but keep a sampler at exposed skin. For diffusion models based on mask inpainting, how to decouple the mask with original clothes also deserves exploring.
@@
Natural clothing deformation. Transformation under intrinsic rules such as STN and TPS is limited for flexible deformation. For example, they cannot put sleeves before the torso. Patch-based methods alleviate the limitation to some extent. Flow estimation learns pixel offset which improves flexibility but cannot generates new pixels that are not displayed in target clothing image but are shown under current posture. Implicit transformation shows potentials in clothing deformation and deserves further exploration, especially driven by a 3D deformation model. Additionally, existing methods distort target clothing image into the try-on area and ignore the original style. For example, a cropped top is stretched or a long coat is shrunken to normal length. Fine-grained clothing landmarks could be helpful for this situation.
@@
Controllable generation with diffusion model. Diffusion models show powerful generalization and high quality, which is appealing for virtual try-on tasks. However, current models only meet semantic-level control such as satisfying the description from text. More precise and strict control has not been well solved such as preserving the original styles (e.g., Fig. 19). Effectively encoding clothing image and injecting to the generation model, or just using diffusion models to refine the warped clothes deserve a try.
@@
Mask-free in-the-wild try-on. Current in-the-wild try-on methods rely on accurate mask for inpainting and only allows slight occlusions. However, fixed masking strategy is not flexible enough and mask inpainting has the following drawbacks for image-based virtual try-on: 1) It is difficult to preserve the original characteristics such as the skin tone, tattoo or other decorations. 2) The original clothing area that is not masked will cause artifacts. 3) The target clothing is distorted into the inpainting area with the loss of clothing types. Therefore, how to accurately learn the clothing area that is expected to be altered is important for mask-free in-the-wild try-on. It also deserves a try to explore flexible loss functions to support unpaired training, which respectively constrains try-on area with target clothing image and non-try-on area from person image.